OpenStudy (anonymous):

which of these ordered pairs is the best estimate of the solution of the system y = 3x + 2 and y = -1/4 x + 1? (0.5, 0.5) (-0.3, 1.1) (1.2, 0.8) (0.8, -0.5)

OpenStudy (anonymous):

(-0.3, 1.1)\[\left\{x\to -\frac{4}{13},y\to \frac{14}{13}\right\} \]\[\{x\to -0.307692,y\to 1.07692\} \]